OpenStudy (anonymous):

the length of a rectangle is 10 cm more than the width. the perimeter is 80 cm, find the dimensions

OpenStudy (dtan5457):

Set up with equation. 2(x+10)+2(x)=80 Factor that out to 2x+20+2x=80. That makes it 4x+20=80 Subtract to 4x=60. x=15, so that would be the width. And since the length is 10 more it would be 25. 25+25+15+15=80, the correct perimeter.